### v1

# Photoelectron Circular Dichroism of Aqueous-Phase Alanine

Abstract
Amino acids and other small chiral molecules play key roles in biochemistry. However, in order to understand how these molecules behave in vivo, it is necessary to study them under aqueous-phase conditions. Photoelectron circular dichroism (PECD) has emerged as an extremely sensitive probe of chiral molecules, but its suitability for application to aqueous solutions had not yet been proven. Here, we report on our PECD measurements of aqueous-phase alanine, the simplest chiral amino acid. We demonstrate that the PECD response of alanine in water is different for each of alanine’s carbon atoms, and is sensitive to molecular structure changes (protonation states) related to the solution pH. For C 1s photoionization of alanine’s carboxylic acid group, we report PECD of comparable magnitude to that observed in valence-band photoelectron spectroscopy of gas-phase alanine. We identify key differences between PECD experiments from liquids and gases, discuss how PECD may provide information regarding solution-specific phenomena — for example the nature and chirality of the solvation shell surrounding chiral molecules in water — and highlight liquid-phase PECD as a powerful new tool for the study of aqueous-phase chiral molecules of biological relevance.
keywords:
Photoelectorn Circular Dichroism, Liquid-Jet Photoelectron Spectroscopy, Chirality
[…]
{tocentry}
1 Introduction
Chirality describes a broad class of objects that share the property of being non-superposible with their mirror images. This general property of chirality may be found at all size scales and levels of complexity in nature, from human hands and sea shells down to small molecules. Many of the molecules most essential for life on Earth are chiral, including macromolecules such as proteins and DNA as well as smaller building blocks of life such as amino acids and sugars. All chiral biomolecules have been found to be overwhelmingly present in only a single enantiomeric form in living organisms, despite the mirror-image enantiomeric forms of a chiral molecule exhibiting identical physical and chemical properties in achiral environments (exempting small anticipated energetic differences between enantiomers due to parity-violating effects in the weak interactions) 1. The origins of this asymmetry in biology remain unresolved. In chiral environments, the chemical activity of the different enantiomers of a chiral molecule can differ dramatically. As such, experimental methods capable of distinguishing between enantiomers under biologically relevant conditions are of direct interest for chemical and life sciences. Absorption-based electronic circular dichroism (ECD) describes the asymmetry in absorption of left- or right-handed circularly polarized light (CPL) in the UV-visible range by different enantiomers of a chiral molecule. ECD is widely used as a means to determine the degree of enantiomeric excess for chiral molecules in solution 2. However, it is dependent on the interaction between the electric and magnetic dipoles of the chiral system with the CPL, and correspondingly the intrinsic effect magnitudes are small, generally on the order of 0.01% 3.
In 2001, it was first experimentally demonstrated that one-photon photoionization of chiral molecules by circularly polarized synchrotron radiation may also be used to differentiate between enantiomers 4. This photoelectron circular dichroism (PECD) — predicted theoretically 25 years earlier by Ritchie 5, 6 — manifests as a forward-backward asymmetry in the measured photoelectron flux along an experimental axis defined by the direction of light propagation. Arising purely due to electric-dipole interactions, PECD is much more pronounced than ECD, exhibiting asymmetries on the order of 1% up to a few 10’sfor randomly oriented molecules 7. Alignment of molecules has been demonstrated to further increase PECD effect sizes 8, 9, 10. PECD appears to be a general feature in photoionization of chiral systems by CPL, having been observed in core-level and valence-band photoemission from terpenes 11, 12, 13, 14 as well as from molecular dimers and larger clusters 15, 16, amino acids 17, 18, 19, organometallic complexes 20, and nanoparticles 18. In addition to the traditional strengths of photoelectron spectroscopy, including chemical-state and site specificity, PECD is uniquely sensitive to small differences in molecular electronic structure and is capable of clearly distinguishing between different structural and conformational isomers of a chiral molecule 21, 22. PECD is a powerful method capable of much more than the resolution of enantiomers and the determination of enantiomeric excess. These very appealing analytical capabilities have driven the growing field of laser-based multi-photon PECD since 2012.23, 24, 25
Water is critically important for biochemistry, playing an active role in determining the functionality of amino acids, proteins, and DNA through structure stabilization and mediation of intra- and intermolecular interactions 26. For chiral molecules in solution, interactions with solvent molecules, even achiral solvents such as water, may profoundly influence both the magnitude and the sign of measured chiroptical effects 27. Additionally, it is possible that achiral solvent molecules may themselves adopt chiral arrangements within the first solvation shell around a chiral solute, mimicking to some extent the solute’s chiral structure 28. Understanding the complex interactions between chiral solutes and achiral solvents is clearly a necessary prerequisite for developing deeper insight into the chemical activity of chiral molecules in solution, whether in biochemical or synthetic contexts. Core-level PECD seems uniquely suited for an investigation into the remarkably active role of solvent molecules in determining chiral solute properties. However, PECD of liquid-phase samples has only very recently become experimentally feasible 29, 30.
The development and growth of liquid-jet photoelectron spectroscopy (LJ-PES) over the past few decades has led to the solution of many of the technical obstacles that long prohibited the study of volatile liquids under the vacuum conditions required for PES 31, 32. However, the application of PECD to liquids, and particularly to aqueous solutions, introduces a number of new experimental challenges that are not present for gas-phase studies. PECD, based upon the scattering and interference of departing partial electron waves, exhibits a strong kinetic-energy dependence, with the most pronounced effects manifesting for photoelectrons with kinetic energy of less than approximately 15 eV 11. For photoelectron spectroscopy of condensed-phase systems this poses a clear challenge due to the unavoidable convolution of primary photoelectron spectral features in this kinetic-energy range with the high-intensity low-energy electron background composed of electrons inelastically scattered in the sample bulk. For the cases of liquid water and of bulk-soluble species in aqueous solution, the opening of efficient quasi-elastic scattering channels within the same kinetic-energy range additionally leads to the broadening of photoelectron features and the homogenization of photoelectron angular distributions 33, 34.
Despite these challenges, we have previously demonstrated the viability of liquid-phase PECD measurements for the chiral molecule fenchone as a neat liquid 30. We now report PECD in core-level photoionization of aqueous-phase alanine — the simplest chiral amino acid — at different solution pH values corresponding to the cationic, zwitterionic, and anionic form of the molecule. We summarize the results of our measurements, highlight the key differences in measuring PECD in the aqueous and gas phases, and discuss the potential influence of water solvation on PECD. We outline new opportunities for aqueous-phase PECD going forward, including the relatively straightforward application of PECD to study cationic or anionic chiral molecules; a topic that has only recently begun to be explored 35, 36, 37. More generally, this study constitutes the first core-level PECD investigation of chiral amino acids in any context. Liquid-phase PECD studies of the elementary constituents of life are highly relevant to in vivo biological conditions. We present a first step toward a bottom-up approach to understanding biomolecular complexity through the preparation of simple aqueous solutions, enabling the study of fragile thermolabile chiral species such as amino acids, peptides, sugars, and nucleic acids, whose vaporization into the gas phase is extremely challenging.
1.1 Experimental Considerations
In this work, we focus on core-level C 1s measurements of aqueous-phase alanine. Alanine has three chemically distinct carbon atoms — a carboxylic acid group, a chiral central carbon adjacent to an amine group, and a methyl group, which we denote C1, C2, and C3, respectively — resulting in a solution-phase C 1s PE spectrum with three distinct features that overlap to varying degrees depending on the solution pH / alanine protonation state (Fig. 1). An investigation of the precisely energy-referenced C 1s spectrum of alanine is not the focus of this work, and we have aligned all spectra discussed below to the C3 peak position (corresponding the alanine’s methyl group, see gray dashed line in Fig. 1) in order to better illustrate the kinetic-energy (KE) shifts occurring for C1 and C2 upon protonation or deprotonation. As we did not employ the energy-referencing scheme outlined in our earlier work 38, 39, 40, the KE axes presented in this work represent as-measured values. These values may be subject to inaccuracies due to spectrometer offsets, the presence of streaming potentials and other stray fields, and changes in the bulk electronic structure of water upon the addition of large amounts of NaOH and HCl needed for pH adjustment 41, 42, 43.
To track chemical shifts upon protonation / deprotonation of alanine’s functional groups, we first measured PE spectra from aqueous-phase alanine for solution pH values of 1, 7, and 13 using 335 eV photons, thereby generating photoelectrons with approximately 40–45 eV KE. These pH values were chosen based on the pKa values of alanine’s functional groups (2.3 and 9.9 for alanine’s carboxylic acid and amine groups in water, respectively) and ensured that, for a given solution pH, >95% (>99% for pH 7 and 13) of the alanine molecules in solution adopted a single charge state. At these rather high KEs the primary PE features are well separated from the high-intensity background region mentioned above, which becomes significant below approximately 20 eV KE.
Under strongly acidic conditions, alanine is predominantly in a cationic state ($Ala^{+}$), with the amine functional group protonated (Fig. 1, top). In this state, despite some overlap between C2 and C3, the three carbon PE features are easily resolvable owing to pronounced chemical shifts arising due to the different chemical environments of the carbon sites. The lowest KE feature corresponds to photoelectrons originating from the relatively electronegative COOH group (C1). The central PE feature corresponds to the chiral center adjacent to the amine group (C2), and the final feature is attributable to the methyl group (C3). These assignments are consistent with previous solid-state and gas-phase work 44, 45. As can be seen by comparing the pH 1 and pH 7 data in Figure 1, deprotonation of the carboxylic acid for zwitterionic Alazw is accompanied by a noticeable chemical shift of the C1 peak (relative to the C3 peak) of approximately 0.71 eV toward higher KE (lower binding energy), reflecting the increasing electron density at that site upon deprotonation. In contrast, C2 does not experience much of a chemical shift, suggesting that the central carbon is not particularly sensitive to the changing electron density on the neighboring functional group. This may be a consequence of effective charge screening by nearby water molecules.46 At this pH, all three primary PE features are still easily resolvable. At pH 13, alanine is found in the anionic form, $Ala^{-}$, with the carboxylic acid deprotonated and the amine group neutral. In this case, we observe a chemical shift of 0.44 eV toward lower KE for the C2 peak, corresponding to an increase in electron density at that group compared to the pH 7 case. This chemical shift results in substantial overlap of the C2 and C3 peaks. We note that the C1 peak remains clearly isolated from the C2 and C3 features for all charge states, thereby simplifying the determination of PECD for this carbon site.
Figure 1: Representative as-measured C 1s photoelectron spectra corresponding to aqueous-phase alanine’s three chemically distinct carbons measured with $h\nu$ = 335 eV photons under acidic (top), neutral (middle), and basic (bottom) aqueous conditions. Measurements were conducted with 1 M aqueous solutions of L-alanine. The dashed black curves are fits to the data following the subtraction of background signals (dashed purple lines) using Gaussian profiles, with the widths of the overlapping C2 and C3 peaks constrained to differ by a maximum of 10%. All spectra were shifted to the as-measured kinetic energy of the C3 peak, corresponding to the methyl group (dashed gray vertical line). Molecular structures of the dominant protonation state of alanine at the various pH values investigated are shown as insets.
Similarly to our previous study of PECD from neat liquid fenchone 30, we determined PECD using synchrotron radiation and an in-vacuum liquid-jet photoemission spectrometer via sequential measurements of photoelectron flux upon photoionization of aqueous alanine solutions with left- and right-handed CPL (see Experimental section below for additional details). The implemented beamline settings corresponding to each type of polarization were confirmed previously through measurements of PECD from gaseous fenchone 29, 30. Electrons were detected in the backward direction, with our electron analyzer mounted at $\approx$50 with respect to the axis of light propagation, approximating a magic-angle geometry of 54.74 29. This angle was chosen due to technical considerations of compatibility with the open port at the P04 soft x-ray beamline, where these measurements were performed.
For single-photon photoionization of a randomly oriented sample 47, the normalized photoelectron angular distribution (PAD) within the electric-dipole approximation is given by:
$I^{p}(\theta)=1+b^{p}_{1}P_{1}(cos\theta)+b^{p}_{2}P_{2}(cos\theta)$ (1)
Here, $I^{p}(\theta)$ is the photoionization intensity at an angle $\theta$ measured with respect to the photon-propagation direction for a given light polarization ($p$ = ±1 for circularly polarized light and 0 for linearly polarized light, respectively). $P_{n}$ represents the Legendre polynomial of the nth order. The two coefficients $b^{p}_{1}$ and $b^{p}_{2}$ completely encompass the target-specific contribution to the PAD under a fixed set of experimental conditions. For achiral molecules, $b^{p}_{1}=0$ and thus the PAD is described by $b^{p}_{2}$, more commonly denoted $\beta$ for experiments involving linearly polarized light (i.e. $b^{0}_{2}=\beta$, while $b^{\pm 1}_{2}=-\beta/2$) 48. For chiral molecules, $b^{\pm 1}_{1}$ may be non-zero and $b^{+1}_{1}=-b^{-1}_{1}$, changing sign upon inversion of the handedness of CPL (or of the enantiomer). An asymmetry factor, $G$, may be defined as the simple difference between left- and right-handed PE intensities divided by their average. Using Eq. 1 and noting that $P_{1}(x)=x$, we then have:
$G=\frac{I^{+1}(\theta)-I^{-1}(\theta)}{[I^{+1}(\theta)+I^{-1}(\theta)]/2}=\frac{2b^{+1}_{1}cos\theta}{1+b^{\pm 1}_{2}P_{2}(cos\theta)}$ (2)
It can be seen that this provides a means to extract $b^{+1}_{1}$ although, in general, a knowledge of $b^{\pm 1}_{2}$ is also necessary. However, $P_{2}(cos(54.74^{\circ}))=0$. Thus, at magic angle:
$b^{+1}_{1}=\pm\frac{[I^{+1}-I^{-1}]}{[I^{+1}+I^{-1}]}\cdot\frac{1}{cos(54.74^{\circ})},$ (3)
the negative sign applying when, as here, measurement is made in the backwards hemisphere. We note that $P_{2}(cos(50^{\circ}))\neq 0$, and we discuss the relevant consequences of our near magic-angle measurement geometry in greater detail below.
As $b^{+1}_{1}$ simply changes sign upon exchange of enantiomer, in principle only a single pair of measurements for a single enantiomer with left- and right-handed CPL is necessary for its determination at any given energy using Eq. 2. The normalized intensities $I^{p}(\theta)$ are obtained from the experimental photoelectron intensities at a specified electron energy and detection angle $\theta$, but require correction accounting for possible variation in photon flux and other sources of medium term experimental drift. Despite our best efforts to ensure that the fluxes of left- and right-handed CPL were equal — monitored via a photodiode — small persistent differences in light intensity of 1–3% remained and so data normalization was essential. In gas-phase PECD studies, $4\pi$ imaging spectrometers enable collection of the total ionization yield, and thus enable direct normalization.12 In the present experimental arrangement the total ionization yield cannot be monitored and an alternative intensity-correction procedure utilizing the baseline signal was adopted. It is assumed that the background of scattered electrons is achiral and essentially independent of light polarization state. We therefore scaled all left- and right-handed CPL data (hereafter denoted as $+$ and -- spectra) to ensure equal intensity of the spectra at the low- and high-KE edges, where only an achiral background is present.
Figure 2: Top: Illustrative aqueous-phase (pH 13) alanine C 1s photoelectron spectra corresponding to right- and left-handed circularly polarized light (red and blue lines, respectively). The high-energy spectra (right) exhibit low background signal, but vanishing photoelectron circular dichroism (PECD). The low-energy spectra sit atop a large background signal, but reveal significant PECD following subtraction of background signal (purple lines). Bottom: Hypothetical $b^{+1}_{1}$ parameters for each of alanine’s carbon groups as a function of peak kinetic energy. The blue shaded zone represents the onset of significant kinetic-energy dependent elastic or quasi-elastic electron scattering. Within the green region, scattering of primary photoelectrons is sufficient to make resolution of the photoelectron features unfeasible. Values of $b^{+1}_{1}$ are currently inaccessible for this system within this kinetic-energy range. Dashed gray boxes represent approximate peak positions for the spectra shown above. The carbon atoms C1, C2, and C3 are identified in Fig. 1.
It is also vital that the baseline is accurately modeled such that the background beneath the peaks of interest can be reliably removed from the intensities $I$ used to determine $b^{+1}_{1}$. It can be seen from Eq. 2 that while any residual background present in the suitably scaled $I^{+}$ and $I^{-}$ will self-cancel in the numerator, this does not apply to the denominator, $[I^{+}+I^{-}]$, whose magnitude would hence be overestimated, and lead to an erroneous underestimation of the derived $|b^{+1}_{1}|$ (see Fig. 3). This challenge is exacerbated for experiments with dilute chiral molecules where the signal-to-background ratio is much reduced (see Fig. 2, top). Details regarding the background-subtraction method employed may be found in the supporting information.
Figure 3: Top: Representative intensity-scaled C 1s photoelectron spectra of 1 M L-alanine aqueous solution (pH 13) measured at $h\nu=305$ eV for left- and right-handed circularly polarized light (blue and red curves, respectively), along with the calculated percent difference between the spectra for C1 group to background subtraction (right axis). Middle: Background-subtracted data, along with calculated percent difference between the spectra for C1 (right axis). Bottom: Values of $b^{+1}_{1}$ obtained through the same process for D-, L-, and racemic DL-alanine (red, blue, and green points, respectively). The error bars of these points represent the standard deviation of the percent-difference data shown in the middle panel. The spectra shown here are averages of ten acquisitions, as typical for the data discussed in this study.
At the near magic-angle geometry of $\theta=50^{\circ}$ (in the backward direction) adopted for these experiments, the polynomial $P_{2}(\cos\theta)\neq 0$ and some residual influence of the $b^{\pm 1}_{2}$ parameter is unavoidably included in the measurement (see Eq. 2). However, there are currently no published values of $b^{\pm 1}_{2}$ for core-level photoionization of alanine. To obtain an estimate of the significance of the remnant $b^{\pm 1}_{2}$ contribution, we calculated the value of $b^{0}_{2}$ ($\beta$) for low-energy (KE = 0 to 25 eV) photoelectrons generated upon C 1s photoionization of gas-phase $Ala^{0}$ (details may be found in the supporting information, see Fig. LABEL:SI-fig:S2). We found that for photoionization of C1, the absolute value of $\beta$ is less than 0.8 within this entire KE range, decreasing to less than 0.5 for photoelectrons with KE < 15 eV. Although aqueous-phase alanine will undoubtedly yield different values of $\beta$ due to differences in conformation and the charge state of its functional groups, this calculation nevertheless provides a rough estimate. As a comparison, the measured value of $\beta$ for O 1s photoionization of gas-phase water at 25 eV KE is approximately 1.6, whereas for liquid water at the same energy it is 0.6 34. For both environments, $\beta$ trends strongly toward 0 for lower photoelectron kinetic energies. This decrease in $\beta$ as measured from liquid versus gaseous water was attributed primarily to highly efficient elastic or quasi-elastic scattering of photoelectrons by water molecules in the liquid. As water remains by far the likeliest scatterer of photoelectrons for the 1 M aqueous solutions of alanine studied in this work, it seems reasonable to expect a similar degree of PAD isotropization for $\beta$ as measured in C 1s photoionization of these solutions. Conservatively assuming a $\beta$ value of 0.5 as an upper limit for aqueous-phase alanine and using Eq. 2 while recalling that $b^{\pm 1}_{2}=-\beta/2$ and simply averaging over the acceptance angle of the analyzer (±$\approx\pm 15^{\circ}$) 29, the effect would be to scale our measured values of $b^{+1}_{1}$ by a factor of 0.97. Such a change does not alter the results of this study, and we do not consider it further in subsequent discussion.
For a given aqueous alanine solution (enantiomer, pH), we conducted C 1s LJ-PES measurements as a function of photon energy in the range of 302–310 eV, resulting in C1 peaks in the KE range of 9–17 eV. For each photon energy, we measured 6–10 PE spectra before changing the polarization of the light, resulting in a measurement time of roughly 30–60 minutes per sample. We compared repeats using the same handedness of CPL for a given photon energy over time to ensure that our experimental conditions were reproducible during the measurement period.
In practice, and particularly due to the pioneering nature of this study, it is advantageous to determine $b^{\pm 1}_{1}$ for both enantiomers in order to detect any influence of instrumental and experimental asymmetry. Confirming the anticipated enantiomeric mirroring of the $b^{+1}_{1}$ parameter assures the molecular origin of the observed asymmetry. Wherever possible, we also made the same observations for solutions of racemic DL-alanine, for which no asymmetry is expected. Due to the time-intensive nature of the experiments, which spanned multiple months of synchrotron measurements across four years, it was necessary to find a balance between the photoelectron KE range and enantiomer and protonation state of alanine investigated. This balance is reflected in the greater number of experimental data points measured for low-KE photoelectrons and for Ala- (see for reference Figs. LABEL:SI-fig:S3, LABEL:SI-fig:S7, and LABEL:SI-fig:S11).
2 Results and discussion
Following the process outlined in the supporting information and using Eq. 2, we determined a single value of $b^{+1}_{1}$ per carbon center for each pair of $+$ and -- measurements for a given enantiomer, solution pH, and photon energy (Fig. 3). The data were then averaged within 250 meV KE windows for visualization. The complete datasets without averaging are presented in the supporting information (Figs. LABEL:SI-fig:S4, LABEL:SI-fig:S8, LABEL:SI-fig:S12). The analysis was carried out blind with respect to the enantiomer, such that we analyzed all the data without knowing whether any given pair of measurements corresponded to a solution of L-, D-, or DL-alanine. An overview of the $b^{+1}_{1}$ parameters obtained for C1 at different pH values is presented in Fig. 4. Beginning with the pH 1 data, corresponding to $Ala^{+}$, weak enantiomer-dependent asymmetry in $b^{+1}_{1}$ may be seen for photoelectrons with KE 9 eV. At this energy, L-alanine exhibit a negative value of $b^{+1}_{1}$, while D-alanine yields a somewhat positive value. These values appear to rapidly trend toward zero at higher kinetic energies, and no significant PECD effect is found above 10 eV. For $Ala^{zw}$ at pH 7, the situation is different. Here, $b^{+1}_{1}$ is close to zero in the KE range of 9–11 eV, but appears to increase in magnitude briefly around 12–13 eV. From 14–16 eV, the data is noisy and does not indicate clear PECD. At 17 eV, $b^{+1}_{1}$ increases once more. The sign of $b^{+1}_{1}$ is the same for the corresponding enantiomer at pH 1. The clearest PECD can be seen for the case of $Ala^{-}$. For this protonation state, $b^{+1}_{1}$ as measured for C1 is clearly non-zero for the entire KE range measured, even up to 17 eV. The sign of $b^{+1}_{1}$ remains constant throughout the entire range, and is consistent with that observed for $Ala^{zw}$ and $Ala^{+}$.
Figure 4: Values of the $b_{1}^{+1}$ photoionization parameter obtained for C 1s measurements of aqueous solutions of D-, L-, and DL-alanine (red, blue, and green points, respectively) at pH 1, 7, and 13 (top, middle, and bottom; corresponding to the cationic, zwitterionic, and anionic form of the molecule, respectively). All $b_{1}^{+1}$ values shown correspond to photoionization of the C1 carboxylic group. The data displayed is the result of binning the results of individual measurement sets using a kinetic-energy window of 250 meV. The shaded error bars represent the combined error of the binned data points. A description of the error propagation may be found in the supporting information.
We were unable to conclusively identify signatures of PECD for the C2 and C3 atoms, with the possible exception of $Ala^{zw}$ in the KE range of 15–19 eV (Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S8). However, due to the limited quantity of data in this KE range for $Ala^{zw}$ and the overlapping nature of these peaks, we do not consider the C2 and C3 data to convincingly demonstrate PECD. As mentioned above, we dedicated the most time to collecting data points for $Ala^{-}$, which appeared from the onset to yield the most clear PECD. As such, we are the most confident in our results for $Ala^{-}.$ Although it may seem surprising that C1, C2, and C3 should exhibit such large differences in PECD magnitude, significant site-dependent PECD has previously been reported in C 1s studies of carvone and fenchone,49, 14 thereby highlighting the sensitivity of PECD to the atomic site probed. The datasets corresponding to Fig.4 for C2 and C3 are available in the supporting information (see Figs. LABEL:SI-fig:S7 and LABEL:SI-fig:S11).
Taken altogether, the data reveal a significant PECD effect for aqueous-phase alanine in photoionization of the C1 carbon group, for all three of alanine’s protonation states. The sign of $b^{+1}_{1}$ is consistent between $Ala^{+}$, $Ala^{zw}$, and $Ala^{-}$, but differences in magnitude are apparent, with the most prominent effects measured for $Ala^{-}$. In contrast to most other PECD studies, we do not note a clear KE dependence for $b^{+1}_{1}$ in our data. This is likely due to the compensation of increasing PECD by concomitantly increasing elastic scattering at low KEs (as illustrated in Fig. 2 within the blue and green regions). The extent to which the measured $b^{+1}_{1}$ values will be reduced by scattering will depend on the specific KE-dependent elastic-scattering cross section. As such, substantial electron scattering may result in changes not only to the magnitude of the measured $b^{+1}_{1}$ values, but also their apparent KE dependency. This result suggests that, for the current case of alanine in water, there may be limited experimental benefit (in terms of magnitude of measured $b^{+1}_{1}$) in focusing on very-low KE photoelectrons. As these low-KE spectra are the most difficult to analyze due to the necessity for significant background subtraction, follow-up studies may benefit from investigation of higher-energy (KE = 15–25 eV) photoelectrons. We can estimate the average degree of $b^{+1}_{1}$ attenuation in our present measurements through comparison with previous work. As mentioned above, the angular asymmetry of the O 1s PAD is reduced by 60% for liquid compared to gaseous water at KE = 25 eV 34. Similarly, in our previous core-level PECD study of liquid fenchone 30, we reported a factor of 5 decrease for $b^{+1}_{1}$ compared to that measured for gas-phase fenchone within the KE range of 9–13 eV 14. This seems to be a general trend upon condensation of chiral molecules; the magnitude of PECD measured for free serine in the gas phase was also found to decrease upon condensation of serine into nanoparticles, although in this case condensation was also accompanied by changes of molecular charge state and conformer population, both of which influence PECD in a non-trivial manner 18. A more detailed understanding of elastic scattering of low-energy photoelectrons in aqueous solutions of alanine will be required to more accurately characterize the degree of isotropization in the present study. However, based on the work mentioned above, we estimate that our present results for $b^{+1}_{1}$ may be attenuated by a factor of 3–5. For the C1 data shown in Fig. 4, this means that our measured values of $b^{+1}_{1}$approx 0.02 likely correspond to initial $b^{+1}_{1}$ values of approximately 0.06–0.1 prior to deleterious photoelectron scattering. The above arguments equally hold for the C2 and C3 carbons and would serve to make what may have been initially low values $b^{+1}_{1}$ for those atoms undetectable with the current experimental sensitivity.
Although no other core-level PECD studies of alanine have been published, we may compare our results with the existing theoretical and experimental investigations of valence-band PECD of alanine. These studies have revealed only modest PECD, with $b^{+1}_{1}~{}\approx~{}0.02$, corresponding to a PECD effect of 4% 17, 50. This value is roughly 10 times larger than the comparable ECD magnitude (g factor) for alanine in the gas phase 51, but significantly lower than the asymmetry reported for valence-band photoemission from rigid molecules such as terpenes ($b^{+1}_{1}$ up to $\approx$ 0.1)52. The relatively weak PECD reported for alanine is similar to that observed for other flexible molecules in the gas phase, including alaninol 53, indanol 54, 37, 22, endoborneol 55, and proline 19, and may have to do with conformational averaging. Reflecting this conformational sensitivity, most PECD studies in the literature focus on relatively rigid molecules, including bicyclic terpenes and terpenoids 4, 11, 56, 12, 13, 15, 52, 57, 30, or other heterocylic oxirane derivatives 58, 59, 60, 8, 61, 62, 10.
If core-level photoionization of alanine’s C2 and C3 atoms in the gas phase were to result in similar values as measured for valence-band photoionization of alanine ($b^{+1}_{1}$approx 0.02). they would likely manifest in our experiments as $b^{+1}_{1}$approx 0.004-0.007 (based on our scattering estimate above). As can be seen in Fig. 4, the conclusive detection of such low values is currently not feasible. Considering the general trend of flexible molecules exhibiting weaker PECD, it appears reasonable that we currently lack the sensitivity to PECD from all of alanine’s carbon groups. Somewhat more surprising is that, despite elastic scattering, our measured values of $b^{1}_{1}$ for core-level photoionization of C1 in aqueous $Ala^{-}$ are comparable to those reported for valence-band photoionization of gas-phase $Ala^{0}$.
The specific details, signs, and magnitudes of PECD arise as results of extensive interference between multiple continuum partial electron waves upon photoionization, and thus defy intuitive prediction 47. Nevertheless, it is clear that the local scattering environment must necessarily influence the continuum wave function of each partial electron wave. The significant role of final-state scattering in core-level PECD was first experimentally demonstrated in C 1s photoionization of gaseous carvone and fenchone, for which strong PECD was observed despite the spherical symmetry of the initial states 49, 14. As our present study constitutes the only report of PECD of aqueous-phase systems at this time, it is useful to briefly discuss how some of the most prominent differences between aqueous-phase and gas-phase molecules may manifest in PECD measurements. Theoretical approaches to determining PECD are currently limited to small gas-phase molecules, and we hope that the following discussion will stimulate the development of new methods capable of modeling PECD for aqueous-phase systems, which we anticipate will prove invaluable for the the application of liquid-jet PECD as a practical analytical technique. We suggest that a model of PECD for zwitterionic alanine microsolvated by 5 explicit water molecules — the minimum required to stabilize the charged forms of alanine in solution 63 — would constitute a valuable step in this direction.
[…]
The most energetically favorable conformations of alanine in the gas phase are stabilized by intramolecular hydrogen-bond formation between the NH2 and COOH groups 65, 66, 67. These intramolecular hydrogen bonds will likely be disrupted in solution, being replaced by intermolecular alanine-solvent interactions. The loss of intramolecular hydrogen bonds may result in subtle changes to PECD, as intramolecular hydrogen-bond orbital polarization has been shown to play a small yet significant role in determining chiral PAD asymmetry 68. The relevance of intermolecular interactions for PECD has been revealed through studies of homochiral clusters of small, weakly interacting molecules, for which profound changes in PECD were observed between isolated monomers, dimers, and larger "n-mers" 69, 16. In the present context, the formation of new intermolecular hydrogen bonds between water and alanine will almost certainly substantially influence the measured PECD. We also note that a molecule’s lowest-energy conformers well may differ between the gas and solution phases 46, and that such differences would be associated with changes in the nuclear scattering potential and the consequent PECD signal.
Although only little work focused on the solvation environment of alanine in solution across its different protonation states exists, we will attempt to sketch a preliminary model based on the current state of the literature. In the case of $Ala^{zw}$, water molecules readily hydrogen bond to the charged COO- and NH+${}_{3}^{+}$ groups, forming a bridge between them and thus stabilizing the doubly ionized form of the molecule 70, 71, 63, 72. At this pH, water molecules slightly prefer bonding to COO-, forming on average six bonds with that group compared to three with NH+${}_{3}^{+}$ 73. This preference of water to form hydrogen bonds with the COO- group is consistent for $Ala^{-}$ and is also similar to what has been previously reported for zwitterionic glycine in water 74. Under acidic conditions, the trend seems to invert, with four waters preferentially coordinating to the NH+${}_{3}^{+}$ group and only two to the COOH group 73. Simulations of alanine in the presence of larger numbers of explicit water molecules have demonstrated that alanine–water complexes are dynamic, with alanine’s different functional groups capable of rotating, and with hydrogen-bonded water molecules within the first solvation shell constantly exchanging with those from the second 70. However, these processes occur on the ps scale, and while they may be significant for time-resolved PECD studies, alanine’s solvation shell appears relatively structured and rigid averaged over the time scale of minutes, with the charged functional groups hydrogen-bonded to water molecules more than 95% of the time 70, 71. Even alanine’s methyl group, which does not form strong hydrogen bonds with water, seems to be involved in the formation of a structured screening water network 75 These studies identify large differences in the degree of hydrogen bonding between water molecules and alanine’s different functional groups depending on pH, which may partially explain the site- and protonation-state-dependent PECD that we have measured.
PECD has recently been reported for gas-phase photoionization of an achiral chromophore weakly bound to a chiral molecule, indicating that PECD exhibits long-range (on the 5 Å scale) sensitivity to chirality 76. As such, PECD of aqueous-phase chiral molecules is almost certain to be sensitive to induced solvent chirality within the first solvation shell. The extent to which induced chirality may arise in a solvation shell appears to depend strongly on solvent polarizability and the van der Waals volume of the solvent molecule 28, 77, 78. Chiral sum frequency generation (SFG) spectroscopy experiments probing induced chirality in the solvation shell around small proteins near the liquid-vapor interface have revealed that the most significant chiral signal from the solvent originates within the first solvation shell, with the second and third solvation shells being nearly achiral 79. These conclusions are in agreement with results from molecular dynamics simulations 78 and suggest an analogy of molecules in the first solvation behaving cooperatively as an achiral glove (for example, a non-handed nitrile glove), adopting a preferential handedness only upon being worn (or upon solvating a chiral solute). The consequences of induced chirality for chiral spectroscopy can be significant, as in the case of methyloxirane in benzene, for which the ECD signal appears to be dominated by the solvation-shell structure rather than by the intrinsically chiral solute itself 80. Vibrational circular dichroism (VCD), probing the difference in absorption of left- and right-handed CPL in the infrared and near-infrared region, has also proven capable of resolving signatures of chirality transfer between solute and solvent via hydrogen bonding 81, 82, 83, 84. Although the chirality of alanine’s aqueous solvation shell has not yet been studied directly, all criteria seem to be met for the formation of a chiral solvation shell. We note that the ECD response of $Ala^{zw}$ differs substantially in the solid vs. in the liquid phase 85 and in general the ECD spectra of aqueous-phase alanine is dependent on solution pH 3, but it is presently unclear whether these trends are related to solvent effects, or simply due to differences in solute conformation or protonation state. We suggest that molecular dynamics studies investigating the chirality of aqueous-phase alanine’s solvation shell as a function of protonation state, combined with new approaches to calculating PECD for larger systems, will be of direct relevance for interpreting our present experimental results. The chirality of a solvation shell will likely manifest differently for different chiral spectroscopic techniques. ECD and VCD appear capable of revealing general chiral effects in the solvation shell, and chiral SFG is well suited for studying molecules at the liquid-vapor interface. Due to its site specificity, core-level PECD may enable much more localized probing of chiral solvation effects, both for bulk-soluble and surface-active molecules. This study constitutes an important first step in this direction.
Importantly, we note that although PECD of aqueous-phase chiral molecules currently requires time-intensive measurement, new experimental approaches that enable simultaneous detection of photoelectrons in both the forward and backward directions with enhanced detection efficiency would dramatically expand the general applicability of this method. In particular, we highlight liquid-compatible velocity map imaging as a technique of considerable interest for the further development of LJ-PECD 86.
3 Conclusions
We have measured core-level photoelectron circular dichroism (PECD) for aqueous-phase alanine, the smallest chiral amino acid, for each of its three charge states (cationic, zwitterionic, and anionic). We detected significant PECD for photoionization of alanine’s carboxylic acid functional group, which exhibited a clear magnitude dependence across pH conditions, but did not exhibit a strong kinetic-energy dependence. We attribute the latter observation to significant kinetic-energy dependent elastic scattering of low-energy primary photoelectrons. We did not measure significant PECD in photoionization of alanine’s chiral center or methyl group, suggesting that the effects for those groups are weak and below our currently modest experimental sensitivity. For photoionization of the carboxylic acid, PECD is the most pronounced under basic solution conditions, where alanine adopts the anionic form, decreases for the zwitterionic form, and is very weak for the cationic form. The differences observed across protonation states may be due to a number of factors, including changes in molecular charge state, functional-group dependent solute-solvent interactions, and first solvation-shell structure. Advances in computational approaches for modeling PECD in multiple-molecule systems will be necessary to identify the relative weights of these effects, and higher collection efficiency detectors capable of simultaneous measurement in the forward and backward direction will dramatically simplify its application in the future. Our demonstration of aqueous-phase PECD marks a significant advance in the field of liquid-jet photoelectron spectroscopy. Given its chemical-state and site specificity and enantiomeric sensitivity, liquid-phase PECD is likely to become a remarkably useful analytical method for studying chiral molecules under biologically relevant aqueous conditions.
4 Experimental
We carried out all aqueous-phase PECD experiments using our custom LJ-PES apparatus — Electronic Structure from Aqueous Solutions and Interfaces (EASI) — developed specifically for this purpose 29. We introduced alanine aqueous solutions into the experimental chamber using a custom quartz capillary, with an inner diameter on the order of 20–30 $\mu$m. Solutions were driven by a high-performance liquid chromatography pump (Shimadzu LC–20AD) at a flow rate of 0.8–1.5 mL/minute, resulting in solution velocities of 20–80 m/s. The solutions were prepared by diluting alanine in water, adjusting solution pH where necessary with concentrated HCl or solid NaOH, and finally diluting the solution to reach a final concentration of 1 M alanine. The solutions were injected horizontally and frozen out after the interaction region via contact with a liquid-nitrogen-cooled cold trap. All PES experiments were performed at the P04 soft x-ray beamline of the PETRA III storage ring at DESY in Hamburg, Germany. Critically, the P04 beamline is equipped with an APPLE-II undulator, enabling experiments with circularly polarized light within the experimentally relevant photon-energy range of 290-320 eV 87. Although not available during our early measurements, recent advancements at the P04 beamline now enable the estimation of the degree of circular polarization for a given set of undulator conditions. Retroactive estimates revealed an expected circular polarization of >99% for all experimental conditions employed in this study. D-, L-, and DL-alanine (>98% purity) were purchased from Sigma-Aldrich and Carl Roth and used as received. Enantiomeric purity was confirmed via intermittent testing using ECD (see Figure LABEL:SI-fig:S1). For additional experimental details regarding our instrumentation or LJ-PES more generally, we direct interested readers to our recent technical publication 29. All data reduction was performed using the Igor Pro analysis software (Wavemetrics,version 9).
{acknowledgement}
We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association (HGF), for the provision of experimental facilities. Parts of this research were carried out at PETRA III at the P04 soft x-ray beamline. Beamtime was allocated for proposals II-20180012, I-20200682, II-20210015, I-20211126, and I-20230378. We thank the P04 beamline staff, in particular Dr. Moritz Hoesch and Jörn Seltmann, for their long-standing support for this project. We also thank the European Molecular Biology Laboratory for access to the circular dichroism spectrophotometer. We thank Prof. Petr Slavíček, Prof. Philipp Demekhin, former group members Dr. Anne Stephansen, Dr. Claudia Kolbeck, Dr. Marvin Pohl, Dr. Karen Mudryk, and Dr. Tillmann Buttersack, as well as current group members Harmanjot Kaur, Qi Zhou, Dr. Lukáš Tomaník, Dr. Nicolas Velasquez, Henrik Haak, and Sebastian Kray for valuable discussion and assistance with the transportation and preparation of our setup (EASI) for measurements at PETRA III.
D.S., U.H., M.P., B.C., and B.W. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. GAP 883759–AQUACHIRAL. F.T. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project 509471550, Emmy Noether Programme. F.T. and B.W. acknowledges support by the MaxWater initiative of the Max-Planck-Gesellschaft. B.W. acknowledges support by the Deutsche Forschungsgemeinschaft (Wi 1327/5-1). S.T. acknowledges support from ISHIZUE 2024 of Kyoto University.
{suppinfo}
The supporting information contains electronic circular dichroism measurements of aqueous solutions of alanine, calculations of the photoionization parameter $b^{0}_{2}$ in the kinetic-energy range of 0–25 eV, a summary of all values of $b^{+1}_{1}$ measured for alanine’s different carbon groups, including values calculated before background subtraction and kinetic-energy averaging was performed, and a description of the means determining the error of the $b^{+1}_{1}$ values, as shown in Figures 3 and 4.
The data of relevance to this study may be accessed at the following DOI: 10.5281/zenodo.13902253.